If it takes one person 15 minutes to paint a room and another person 45 mintues topaint the same room, how long will it take if they work together?
Answer by Cromlix(4381)

One person can paint a room in 15 minutes.
Another person can paint the same room in 45 minutes.
One person's rate = 1/15 job/min
Another person's rate = 1/45 job/min
Together rate = 1/x job/min
1/15 + 1/45 = 1/x
Common denominator = 675x
45x + 15x = 675
60x = 675
x = 11.25 minutes
x = 11 minutes and 15 seconds.
Hope this helps.
:-)